OpenStudy (anonymous):

When you see a traffic light turn red, you apply the brakes until you come to a stop. Suppose your initial speed was 11.2 m/s, and you come to rest in 34.7 m. How much time does this take? Assume constant deceleration.

OpenStudy (jamesj):

Let v be the velocity and a the constant deceleration. Then \[ v = 11.2 - at \] This also means that the displacement s is \[ s = 11.2t - at^2/2 \] You now have two equations. You're also given two pieces of information. Suppose it takes t = T time to stop. Then you know that v = 0 when t = T, \[ 0 = v = 11.2 - at \] You also know that this happens when s = 34.7. Hence you have two equations, two unknowns. Solve them for T and you'll find a as well.

OpenStudy (jamesj):

you there? Make sense?

OpenStudy (jamesj):

To make it even more explicit, the two equations are \[ 0 = 11.2 - aT \] \[ 34.7 = 11.2T - aT^2/2 \]